OpenStudy (anonymous):

A garden has 8 white roses and 11 red roses. One rose is plucked from the garden and then one more rose is plucked, without replacement. Let the event W = plucking a white rose the first time; and event R = plucking a red rose the second time. In two or more complete sentences, explain the meaning of probabilities P(R/W) and P(W/R). Do they have same value?

OpenStudy (anonymous):

\(P(W|R)\) means the probability of W given that R has occurred \(P(R|W)\) means the probability of R given that W has occurred

OpenStudy (anonymous):

replace the letter \(W\) by " plucking a white rose the first time" replace the letter \(R\) by "plucking a red rose the second time. " to get your complete sentences

OpenStudy (anonymous):

okay so R/W equals plucking a red rose the second time/plucking a white rose the first time and W/R equals plucking a white rose the first time/plucking a red rose the second time? @satellite73

OpenStudy (anonymous):

lets see if we can say it in idiomatic english rather than math english that way we can understand it better

OpenStudy (anonymous):

not to be picky but \(R|W\) does not "equal" anything it is an event it is the event that the second rose chosen was red, given that you know the first rose chosen was white

OpenStudy (anonymous):

and W|R is the event that the first rose chosen is white, given that the second rose was red kinda weird, like going backwards in time

OpenStudy (anonymous):

that makes \[P(R|W)\] the probability that the second rose chosen is red, given that the first one was white now it is a number (not an event) you can compute this in your head let me know if you need help doing it

OpenStudy (anonymous):

would P(R/W) be P(.61/.42)? @satellite73

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

you know you picked a white one first how many roses are left, and how many are red?

OpenStudy (anonymous):

there is 18 left and 11 are red

OpenStudy (anonymous):

right (i had to scroll up to check )

OpenStudy (anonymous):

that makes \[P(R|W)\] obvious especially if we understand what it means you basically just answered the question

OpenStudy (anonymous):

so P(R/W) means that there is 18 roses left and 11 of them are red? and P(W/R) means that there are 18 roses left and 7 of them are white?

OpenStudy (anonymous):

btw \( P(R/W) = P(.61/.42)\) makes no sense whatsoever \(R|W\) is an event, not a number perhaps you meant \[P(R|W)=\frac{.61}{.42}\] but that doesn't make sense either it could be \[\frac{61}{42}\] but that is a number larger than one, so it is not a probability lets go slow

OpenStudy (anonymous):

\[P(R|W)\] is a number it is the probability that the second one is red, given that the first one is white you can compute this in your head you already said there are 18 roses left and 11 red ones that makes \[P(R|W)=\frac{11}{18}\] just like that !

OpenStudy (anonymous):

Okay so P(W/R) is the probability that the first rose is white P(W/R) = 8/19

OpenStudy (anonymous):

Let the event W = plucking a white rose the first time;

OpenStudy (anonymous):

\[P(W)\] is the probability the first rose is white that is \(\frac{8}{19}\)

OpenStudy (anonymous):

we know \[P(R|W)=\frac{11}{18}\] also at the moment we do not know \[P(W|R)\]

OpenStudy (anonymous):

it is an odd question if you think about it \[P(W|R)\] is the probability that the first rose was white if you know the second one was red like going back in time

OpenStudy (anonymous):

they do not have the same value and the second one is more complicated to compute

OpenStudy (anonymous):

okay well thank you for trying to help me @satellite73

OpenStudy (anonymous):

yw make sure you understand the different between \(P(A)\) which is a number, and \(A\) which is an event we can do the last one if you like, but it is a bit complicated you use something called "baye's formula'

OpenStudy (anonymous):

If you dont mind, i would really appreciate your help with the last one @satellite73

OpenStudy (anonymous):

A garden has 8 white roses and 11 red roses. One rose is plucked from the garden and then one more rose is plucked, without replacement. Let the event W = plucking a white rose the first time; and event R = plucking a red rose the second time. \[P(W|R)=\frac{P(W\cap R)}{P(R)}\] the numerator is easy it is \[P(W\cap R)=\frac{8}{19}\times \frac{11}{18}\]

OpenStudy (anonymous):

the denominator is more complicated , it is \[P(R)=P(R\cap W)+P(R\cap W^c)\] \[=\frac{8}{11}\times \frac{11}{18}+\frac{11}{19}\times \frac{7}{18}\]

OpenStudy (anonymous):

damn typo!!

OpenStudy (anonymous):

\[=\frac{8}{19}\times \frac{11}{18}+\frac{11}{19}\times \frac{7}{18}\]

OpenStudy (anonymous):

final answer \[\frac{\frac{8}{19}\times \frac{11}{18}}{\frac{8}{19}\times \frac{11}{18}+\frac{11}{19}\times\frac{7}{18}}\]

OpenStudy (anonymous):

i will let you compute those numbers i will also let you figure out where they came from gotta run

OpenStudy (anonymous):

okay thank you so much! @satellite73